# ZipAR: Accelerating Auto-regressive Image Generation through Spatial Locality

### v2

@@ 4.1 Implementation Details
To assess the effectiveness of our proposed method, we integrate it with three state-of-the-arts auto-regressive visual generation models: LlamaGen [17], Lumina-mGPT [14] and Emu3-Gen [24]. For text-guided image generation with LlamaGen, we generate 30,000 images and compute CLIP scores against MS-COCO 2014-val dataset with CLIP ViT-B/32 model [15]. For class-conditional image generation with LlamaGen on ImageNet, we report the widely adopted Frechet Inception Distance (FID) to evaluate the performance. We sample 50,000 images and evaluate them with ADM’s TensorFlow evaluation suite [5].
4.2 Quantitative Results
@@ 4.2.1 Class-conditional Image Generation
In this subsection, we quantitatively assess the performance of class-conditional image generation on the ImageNet $256\times 256$ benchmark using the LlamaGen model, as detailed in Table 1. The model employs a $24\times 24$ feature map and necessitates 576 forward passes to generate an image under the next-token prediction (NTP) paradigm. By incorporating ZipAR with a window size of 16, the number of forward passes is reduced by 30.5%, with only a 0.16 increase in FID. Fine-tuning the pre-trained model for a single epoch with an ZipAR-oriented attention mask consistently enhances performance. Specifically, ZipAR-16 reduces the number of forward passes by 30.5% while achieving a lower FID compared to the original model with NTP paradigm. ZipAR-12 further reduces the number of forward passes by 45.8% while maintaining an FID of 3.49.
Table 1: Quantitative evaluation on ImageNet $256\times 256$ benchmark. Here, “NTP” denotes the next-token prediction paradigm. “ZipAR-$n$” denotes the ZipAR paradigm with a window size of $n$. “Step” is the number of model forward passes required to generate an image. The latency is measured with a batch size of 1.
@@
LlamaGen-L NTP 576 15.20 3.16
ZipAR-16 400 (-30.5%) 10.38 (-24.7%) 3.32
ZipAR-12 312 (-45.8%) 8.38 (-39.2%) 4.50
ZipAR-8 224 (-61.1%) 6.56 (-52.4%) 10.22
@@ 4.2.2 Text-guided Image Generation
In this subsection, we quantitatively evaluate the effects of ZipAR on text-guided image generation performance. The LlamaGen-XL model processes a feature map of $32\times 32$ and requires 576 forward passes to generate an image under the next-token prediction (NTP) paradigm. As shown in Table 2, ZipAR-12 results in minimal performance degradation, with a 0.001 decrease in the CLIP score, while reducing inference latency by 61.0%. Notably, ZipAR-4 significantly reduces the inference latency by 83.3% and retains good semantic information, as indicated by a CLIP score of 0.280. Visualization results are presented in Figures 1 and 6. Moreover, we observe that the acceleration ratio for the text-to-image LlamaGen-XL model is higher than that for the class-conditional LlamaGen-L model, primarily due to the higher spatial resolution of the feature maps and generated images. This suggests that ZipAR achieves greater generation efficiency improvements when generating higher-resolution images.
Table 2: Quantitative evaluation on MS-COCO dataset. Here, “NTP” denotes the next-token prediction paradigm. “ZipAR-$n$” denotes the ZipAR paradigm with a window size of $n$. “Step” is the number of model forward passes required to generate an image. The latency is measured with a batch size of 1.
@@
LlamaGen-XL NTP 1024 33.17 0.287
ZipAR-16 544 (-46.8%) 17.65 (-46.8%) 0.287
ZipAR-12 424 (-58.6%) 12.91 (-61.0%) 0.286
ZipAR-8 304 (-70.3%) 9.44 (-71.5%) 0.285
ZipAR-4 184 (-82.0%) 5.51 (-83.3%) 0.280
@@ 4.3 Qualitative Visualizations
In this subsection, we present non-cherry-picked visualizations of images generated using the next-token prediction (NTP) paradigm and the proposed ZipAR framework across Emu3-Gen [24], LlamaGen [17], and Lumina-mGPT [14], as shown in Figures 5 to 7. Notably, ZipAR can reduce the number of model forward steps by up to 91%, 82%, and 67% for Emu3-Gen, LlamaGen, and Lumina-mGPT, respectively, while still producing high-fidelity images rich in semantic information.
Figure 5: Samples generated by Emu3-Gen model with next-token prediction paradigm (the first column) and ZipAR under different configurations (the right three columns). The classifier-free guidance is set to 6.0.
Figure 6: Samples generated by LlamaGen-XL model with next-token prediction paradigm (the first column) and ZipAR under different configurations (the right three columns). The classifier-free guidance is set to 7.5.
Figure 7: Samples generated by Lumina-mGPT-7B-768 model with next-token prediction paradigm (the first column) and ZipAR under different configurations (the right three columns). The classifier-free guidance is set to 3.
5 Conclusion and Future Work
In this paper, we have proposed ZipAR, a new parallel decoding framework designed to accelerate auto-regressive visual generation. ZipAR leverages the spatial locality inherent in visual content and predicts multiple spatially adjacent visual tokens in a single model forward pass, thereby significantly enhancing generation efficiency compared to the traditional next-token-prediction paradigm. Extensive experiments demonstrate that ZipAR can reduce the number of model forward steps by up to $91\%$ on the Emu3-Gen model with minimal impact on image quality.
In the future, we anticipate that integrating ZipAR with other methods that employ the next-set-prediction paradigm, such as Medusa [2] and Jacobi decoding [16], will further enhance acceleration ratios.

### v3

@@ 4.1 Implementation Details
To assess the effectiveness of our proposed method, we integrate it with three state-of-the-arts autoregressive visual generation models: LlamaGen (Sun et al., 2024), Lumina-mGPT (Liu et al., 2024a) and Emu3-Gen (Wang et al., 2024b). All experiments are conducted with Nvidia A100 GPUs and Pytorch framework. For class-conditional image generation with LlamaGen on ImageNet, we report the widely adopted Frechet Inception Distance (FID) to evaluate the performance. We sample 50000 images and evaluate them with ADM’s TensorFlow evaluation suite (Dhariwal & Nichol, 2021).
4.2 Main Results
Table 1: Quantitative evaluation on ImageNet $256\times 256$ benchmark. The generated images are $384\times 384$ and resized to $256\times 256$ for evaluation. Here, “NTP” denotes the next-token prediction paradigm. “ZipAR-$n$” denotes the ZipAR paradigm with a minimal window size of $n$. “Step” is the number of model forward passes required to generate an image. The latency is measured with a batch size of 1.
[…]
LlamaGen-L (cfg=2.0) NTP 576 15.20 3.16
SJD (Teng et al., 2024) 367 (-36.3%) 10.83 (-28.8%) 3.85
ZipAR-16 422 (-26.7%) 11.31 (-25.6%) 3.14
ZipAR-14 378 (-34.4%) 10.16 (-33.2%) 3.44
ZipAR-12 338 (-41.3%) 9.31 (-38.8%) 3.96
LlamaGen-XL (cfg=2.0) NTP 576 22.65 2.83
SJD (Teng et al., 2024) 335 (-41.8%) 13.17 (-41.8%) 3.87
ZipAR-16 423 (-26.6%) 16.46 (-27.3%) 2.87
ZipAR-14 378 (-34.4%) 14.89 (-34.3%) 3.16
ZipAR-12 331 (-41.8%) 13.17 (-41.8%) 3.67
@@ 4.2.1 Class-conditional Image Generation
In this subsection, we quantitatively evaluate the performance of class-conditional image generation on the ImageNet $256\times 256$ benchmark using the LlamaGen model, as summarized in Table 1. The model processes a $24\times 24$ feature map and requires 576 forward passes to generate an image under the next-token prediction (NTP) paradigm. For the LlamaGen-L model, integrating ZipAR with a minimal window size of 16 reduces the number of forward passes by 26.7% without increasing the FID score. For the LlamaGen-XL model, ZipAR-12 achieves a lower FID (3.67 vs. 3.87) while requiring fewer steps than the previous parallel decoding algorithm, SJD (Teng et al., 2024) (331 steps vs. 335 steps). This highlights the efficiency of ZipAR in decoding spatially adjacent tokens in parallel.
Table 2: Quantitative results on diverse automatic evaluation approaches. Here, “NTP” denotes the next-token prediction paradigm. “ZipAR-$n$” denotes the ZipAR paradigm with a minimal window size of $n$. “Step” is the number of model forward passes required to generate an image.
@@
LlamaGen-XL-512 NTP 1024 0.6439 0.2647 -0.0818 5.38
ZipAR-15 562 0.6534 0.2637 -0.0690 5.39
ZipAR-11 451 0.6581 0.2630 -0.0982 5.37
ZipAR-7 324 0.6410 0.2625 -0.1683 5.33
ZipAR-3 185 0.6343 0.2599 -0.3121 5.32
Lumina-mGPT-768 NTP 2352 0.6579 0.2743 0.4164 6.10
ZipAR-20 1063 0.6595 0.2747 0.3971 6.13
ZipAR-17 915 0.6433 0.2732 0.3049 6.12
ZipAR-14 740 0.6589 0.2739 0.3646 6.10
ZipAR-11 588 0.6490 0.2730 0.2861 6.10
Lumina-mGPT-1024 NTP 4160 0.6718 0.2762 0.4232 5.97
ZipAR-20 1331 0.6705 0.2761 0.3913 5.95
ZipAR-17 1150 0.6797 0.2761 0.4018 5.94
ZipAR-14 964 0.6732 0.2747 0.3298 5.94
ZipAR-11 772 0.6723 0.2746 0.3222 5.95
Table 3: Quantitative evaluation on MS-COCO dataset. Here, “NTP” denotes the next-token prediction paradigm. “ZipAR-$n$” denotes the ZipAR paradigm with a minimal window size of $n$. “Step” is the number of model forward passes required to generate an image. The latency is measured with a batch size of 1.
[…]
LlamaGen-XL-512 NTP 1024 33.17 0.287
SJD (Teng et al., 2024) 635 (-38.0%) 24.80 (-25.2%) 0.283
ZipAR-15 562 (-45.1%) 18.98 (-42.7%) 0.287
ZipAR-11 451 (-55.9%) 14.65 (-55.8%) 0.286
ZipAR-7 324 (-68.4%) 10.24 (-69.1%) 0.285
ZipAR-3 185 (-81.9%) 5.86 (-82.3%) 0.281
Luming-mGPT-7B-768 NTP 2352 91.70 0.313
SJD (Teng et al., 2024) 1054 (-55.2%) 60.27 (-34.2%) 0.313
ZipAR-20 1063 (-54.8%) 63.28 (-31.0%) 0.314
ZipAR-17 915 (-61.0%) 58.54 (-36.2%) 0.314
ZipAR-14 740 (-68.5%) 53.41 (-41.8%) 0.313
ZipAR-11 588 (-75.0%) 50.32 (-45.1%) 0.312
@@ 4.2.2 Text-guided Image Generation
In this subsection, we expand our evaluation by assessing ZipAR’s performance using multiple metrics, including VQAScore (Lin et al., 2024), Human Preference Score v2 (HPSv2) (Wu et al., 2023), ImageReward (Xu et al., 2023), and Aesthetic Score, across three models: LlamaGen-XL-512, Lumina-mGPT-768, and Lumina-mGPT-1024, as presented in Table 2. For the LlamaGen-XL model, ZipAR-15 reduces the number of generation steps by 45.1% without any decline in the VQAScore, Image Reward and Aesthetic Score. Similarly, for the Lumina-mGPT-768 model, ZipAR-20 achieves a 54.8% reduction in generation steps while improving VQAScore, HPSv2, and Aesthetic Score. When evaluating the CLIP Score over the LlamaGen-XL model, compared to the previous parallel decoding algorithm, SJD (Teng et al., 2024), ZipAR-7 significantly improves efficiency (324 steps vs. 635 steps) while achieving a higher CLIP score (0.285 vs. 0.283). Moreover, we observe that the acceleration ratio for both text-to-image models is higher than that for the class-conditional LlamaGen-L model. This is primarily attributed to the larger spatial resolution of the feature maps and the generated images. These results suggest that ZipAR provides greater efficiency gains when generating higher-resolution images.
4.3 Ablation Study
4.3.1 Effect of Adaptive Window Size Assignment
In this subsection, we evaluate the effectiveness of the proposed adaptive window size assignment scheme. Specifically, we compare the performance of ZipAR with fixed and adaptive window sizes over class-conditional LlamaGen-L model, respectively. As shown in Figure 6, under similar generation steps, ZipAR with adaptive window size consistently achieves a lower FID than its fixed-window counterpart, which suggests that dynamically adjusting the window size based on token position and context enhances the fidelity of generated images.
Figure 6: Performance comparisons of ZipAR over class-conditional LlamaGen-L model with fixed window size and adaptive window size. Under similar step budget, ZipAR with adaptive window size always achieves lower FID.
4.3.2 Impact on Optimal Sampling Hyperparameters
As presented in Tables 4-5, we performed a grid search to determine the optimal token-sampling hyperparameters, namely, sampling temperature and classifier-free guidance scale, for ZipAR. The results are shown below. These results indicate that ZipAR sampling does not alter the optimal sampling temperature and classifier-free guidance scale.
Table 4: The performance of LlamaGen and ZipAR under different classifier-free guidance. Here, ”*” denotes the results obtained from LlamaGen’s paper.
@@ 4.3.2 Impact on Optimal Sampling Hyperparameters
LlamaGen-L* 1.5 4.74
1.75 3.15
2.0 3.07
2.25 3.62
ZipAR-16 1.5 6.18
1.75 3.72
2.0 3.14
2.25 3.44
Table 5: The performance of LlamaGen and ZipAR under different sampling temperatures. Here, ”*” denotes the results obtained from LlamaGen’s paper.
[…]
LlamaGen-L 0.96 3.53
0.98 3.24
1.0* 3.07
1.02 3.14
ZipAR-16 0.96 3.35
0.98 3.25
1.0 3.14
1.02 3.34
@@ 4.4 Qualitative Visualizations
In this subsection, we present non-cherry-picked visualizations of images generated using the next-token prediction (NTP) paradigm and the proposed ZipAR framework over Emu3-Gen (Wang et al., 2024b) and Lumina-mGPT-7B (Liu et al., 2024a), as shown in Figures 1 and 7. Notably, ZipAR can reduce the number of model forward steps by up to 91% for Emu3-Gen, while still producing high-fidelity images rich in semantic information.
Figure 7: Samples generated by the Lumina-mGPT-7B-768 model with next-token prediction paradigm (the first column) and ZipAR under different configurations (the right three columns). The classifier-free guidance is set to 3.
5 Conclusion
In this paper, we have proposed ZipAR, a new parallel decoding framework designed to accelerate autoregressive visual generation. ZipAR leverages the spatial locality inherent in visual content and predicts multiple spatially adjacent visual tokens in a single model forward pass, thereby significantly enhancing generation efficiency compared to the traditional next-token-prediction paradigm. An adaptive local window assignment scheme with rejection sampling is employed, ensuring that each token is generated with sufficient contextual information. Extensive experiments demonstrate that ZipAR can reduce the number of model forward steps by up to $91\%$ on the Emu3-Gen model with minimal impact on image quality.
In the future, we anticipate that integrating ZipAR with other methods that employ the next-set-prediction paradigm, such as Medusa (Cai et al., 2024) and Jacobi decoding (Santilli et al., 2023), will further enhance acceleration ratios.
Acknowledgements
This work was supported by National Key Research and Development Program of China (2022YFC3602601).
Impact Statement
The proposed ZipAR framework stands out for its high efficiency, which carry significant implications in reducing the carbon emissions attributed to the widespread deployment of deep generative models. However, similar to other deep generative models, ZipAR has the potential to be utilized for producing counterfeit images and videos for malicious purposes.